OpenStudy (anonymous):

Is the ordered pair (3/4,0) a solution of 3x+y>3?

OpenStudy (kropot72):

Plugging the given values into the left hand side gives: \[\large 3x+y=(\frac{3\times3}{4})+0=2\frac{1}{4}\]

OpenStudy (mokeira):

No, it is not a solution Substitute the values and see

OpenStudy (kropot72):

Who thinks the ordered pair is a solution?

OpenStudy (mokeira):

@kropot72 is right.

OpenStudy (midhun.madhu1987):